# The Superstatistics of confined ants: From serenity to panic

### v2

@@ VI Panic response
We now discuss the next 10 minutes after the repellent was injected into the DANGER cell, using data from approximately 50 trajectories of ants in the SAFE arena that were randomly chosen and tracked across all repetitions of the THIN protocol.
Firstly, we found that the mean acceleration maintains the same linear relationship with speed values (Figure S4(left)). A clear time-scale separation is still present, with the slow time scale, $\tau_{2}$, remaining approximately constant over time (Figure 6A). Note that just after a duration of $\tau_{2}$ units, the local kurtosis spreads out, allowing us to assume local equilibrium within the panic regime whenever $t<\tau_{2}$. No significant changes were observed in the persistence time $\gamma^{-1}$ (Figure S4(right)), or in the statistics of the diffusion coefficient, $D$, as indicated by the shape invariance of centered acceleration densities in Figure 6B. The tails of the velocity distributions become stretched, as shown in Figure 6C. This observation aligns with the gradual clustering dynamics under panic conditions, as depicted in the SAFE cell of Figure 2C. Non-gaussianity again arises as a direct combination of a $v$-dependent noise intensity and Superstatistics.
@@
As mentioned previously, the collective behavior of ants in the SAFE cell is mediated by a fast dynamics that span approximately one minute as soon as the repellent is injected. Before that, ants were dispersed throughout the arena. Therefore, when panic is induced in the 5th minute of the experiment, it can be assumed that an ant at the barrier in SAFE is immediately informed and becomes “activated”. After a certain characteristic time, $\tau_{f}$, this ant communicates the danger information to another nestmate, resulting in two activated ants. This cascade continues, and it ends after $T=1\,\mathrm{min}$ when the total number of individuals, $N_{T}=20$, in the arena are informed. Consequently, the number of activated ants at time $t$, $N(t)$, obeys the differential equation $\dot{N}=\lambda_{f}N$ with the initial condition $N(0)=1$, which implies that $\tau_{f}=T/\ln N_{T}=20\,\mathrm{s}$.
@@ VII Conclusions
We have presented controlled laboratory experiments with confined leaf-cutter ants of the species Atta insularis, where ants in a confined space were studied in “normal” conditions for 5 minutes, and in the presence of panic information emerging from the left side of the cell, for another 10 minutes. The experiments show a shift of the center of mass of the ants towards the right as panic is switched on, and the formation of a stable cluster in the last minutes of the experiment.

### v3

@@ VI Panic response
We now discuss the next 10 minutes after the repellent was injected into the danger cell, using data from 48 trajectories of ants in the safe arena that were randomly chosen and tracked across all repetitions of the THIN protocol.
Firstly, we found that the mean acceleration maintains the same linear relationship with speed values (Figure S4(left)). A clear time-scale separation is still present, with the slow time scale, $\tau_{2}$, remaining approximately constant over time (Figure 6A). Note that, after a duration of $\tau_{2}$ units, the local flatness spreads out, indicating that local Gaussians of the form of Equation (2) emerge, allowing us to assume local equilibrium within the panic regime whenever $t<\tau_{2}$. No significant changes were observed in the persistence time $\gamma^{-1}$ (Figure S4(right)), or in the statistics of the diffusion coefficient, $D$, as indicated by the shape invariance of centered acceleration densities in Figure 6B. The tails of the velocity distributions become stretched, as shown in Figure 6C. This observation aligns with the gradual clustering dynamics under panic conditions, as depicted in the safe cell of Figure 2C. Non-gaussianity again arises as a direct combination of a $v$-dependent noise intensity and Superstatistics.
@@
As mentioned previously, the collective behavior of ants in the safe cell is mediated by a fast dynamics, i.e., the blue curve in Figure 2G begins to increase almost immediately. This happens so quickly that the fluctuations obscure the initial rise, but by minute 6, the curve is clearly trending upward. So we estimate the response time to be on the order of one minute. Before that, ants were dispersed throughout the arena. Therefore, when panic is induced in the 5th minute of the experiment, it can be assumed that an ant at the barrier in safe is immediately informed and becomes “activated”. After a certain characteristic time, $\tau_{f}$, this ant communicates the danger information to another nestmate, resulting in two activated ants. This cascade continues, and it ends after $T=1\,\mathrm{min}$ when the total number of individuals, $N_{T}=20$, in the arena are informed. Consequently, the number of activated ants at time $t$, $N(t)$, obeys the differential equation $\dot{N}=\lambda_{f}N$ with the initial condition $N(0)=1$, which implies that $\tau_{f}=T/\ln N_{T}=20\,\mathrm{s}$.
@@
Table 1: Values for model parameters fitted to data.
$\ell_{w}$ $a_{0}$ $\gamma^{-1}$ $\Sigma$ $\Delta\Sigma$ $v_{\sigma}$
0.6 1.0 0.53 0.34 1.99 1.99
$s$ $\tau_{f}$ $\tau_{s}$ $L$ $t_{p}$ $\tilde{x}$
1.27 20 3.2 6.0 5.5 3.8
@@ VII Conclusions
We have presented controlled laboratory experiments with confined leaf-cutter ants of the species Atta insularis, where ants in a confined space were studied in “normal” conditions for 5 minutes, and in the presence of panic information emerging from the left side of the cell, for another 10 minutes. The experiments show a shift of the center of mass of the ants towards the right as panic is switched on, and the formation of a stable cluster after a few minutes.